# Mastering NIM and Impartial Games with Weak Neural Networks: An AlphaZero-inspired Multi-Frame Approach

### v2

@@ 4.3 Multi-frame possibility results
4.4 Two-frame local restoration rule
@@
Suppose a player moves from a position $P$ with $\operatorname{nim}(P)=0$ to a position $Q$ (opponent’s turn). Let the opponent move to $Q^{\prime}$. Then
$\operatorname{nim}(Q^{\prime})\;=\;\Delta(Q,Q^{\prime}).$
If the player can respond with a move from $Q^{\prime}$ to $R$ satisfying
@@
We use $\Delta(A,B)=\operatorname{nim}(A)\oplus\operatorname{nim}(B)$ and XOR associativity:
$\operatorname{nim}(R)\;=\;\operatorname{nim}(Q^{\prime})\oplus\Delta(Q^{\prime},R)\;=\;\Delta(Q,Q^{\prime})\oplus\Delta(Q,Q^{\prime})\;=\;0.$
Proposition 4.6 is structural: it describes a response pattern that keeps returning the game to nim-sum zero, provided the player can (i) compute $\delta=\Delta(Q,Q^{\prime})$ from the last two frames and (ii) choose a legal move $Q^{\prime}\to R$ that realizes the same $\delta$. In NIM, when the opponent changes exactly one heap, $\delta$ is simply $x_{i}\oplus x_{i}^{\prime}$, and testing whether a candidate response move realizes $\delta$ is a local check on the heap being changed.
@@ 4.5 Implementation overview
  • an $\mathrm{AC}^{0}$ selection mechanism for a response move realizing a target $\delta$.
@@ 5 Learning impartial games with neural networks
Our results isolate a sharp distinction:
  • Weak learning: in special structured cases, single-frame agents can succeed via simple symmetries.
  • Strong learning: for general positions, strong mastery requires exploiting history and local invariants (nimber differences) rather than learning global parity.
5.1 Practical considerations
  • What single-frame networks miss. For large boards, fixed-precision bounded-depth networks cannot reliably encode the global nim-sum parity needed for strong mastery (Section 4.2). A single-frame policy/value network therefore has no direct way to base its decisions on the true win/loss structure of arbitrary positions.
  • What two frames add. With two consecutive positions $(Q,Q^{\prime})$, the agent can infer which heap changed and the induced nimber difference $\delta=\Delta(Q,Q^{\prime})$ (Lemma 4.5). Proposition 4.6 shows that conditioning the response move on this locally computable $\delta$ is enough to restore the zero nim-sum invariant after the opponent’s move.
  • Training signals and empirical protocols. Section 6 studies two complementary empirical settings. In Section 6.2 we isolate restoration as a supervised learning task using teacher–random transitions and train policies directly by cross-entropy to the set of optimal restoration moves. We also describe an AlphaZero-style teacher-guided MCTS training protocol (Section 6.3), which uses standard visit-count policy targets from MCTS and value targets from game outcomes. In all cases, simulations and rollouts are ordinary legal play; no special constraints are imposed on rollout moves beyond legality.
  • Coverage and robustness. To avoid collapsing onto a narrow set of deterministic optimal trajectories and to expose the agent to a broad range of transition patterns, we use a near-perfect teacher that plays optimally with high probability but makes a random legal move with small probability $\varepsilon$ (Section 5.2).
5.2 Teacher-guided training as a controlled source of transition data
Let $\pi^{*}$ denote an optimal NIM policy (equivalently, a policy that moves to a zero nim-sum position whenever possible). For a user-specified noise parameter $\varepsilon\in[0,1]$, we define a noisy expert (teacher) policy $\pi_{T}^{\varepsilon}$ as follows: on each move, with probability $1-\varepsilon$ the teacher selects an action according to $\pi^{*}$, and with probability $\varepsilon$ it selects a uniformly random legal move. During training the learner plays against $\pi_{T}^{\varepsilon}$ (and, when desired, uses the same teacher policy inside rollouts).
The occasional random moves serve two roles. First, they increase state-space coverage and prevent the training distribution from collapsing to a small set of optimal trajectories. Second, they generate a richer set of observed transitions $(Q,Q^{\prime})$, which is precisely the information exploited by two-frame representations and the local restoration principle (Proposition 4.6).
5.3 Approximate majority, parity, and why $\mathrm{AC}^{0}$ is the right “feasible” model
A key modelling issue is the following.
The apparent objection.
One could argue that neural networks can compute majority “in practice”, and therefore perhaps $\mathrm{TC}^{0}$-power (which includes majority gates) should be considered feasible. But then $\mathrm{TC}^{0}$ contains $\mathrm{AC}^{0}$, and yet parity remains empirically hard to learn and generalise [20, 21, 6]. How can majority be feasible while parity is not?
Resolution: feasible $\neq$ exact; sharpness matters.
The practical majority phenomenon is best understood as approximate majority: a classifier that gets the label right with high probability under natural noise/typical data distributions, but does not implement a perfectly sharp threshold at exactly $n/2$ for arbitrarily large $n$. Fixed precision and noisy training/inference make perfectly sharp thresholds unstable: tiny perturbations in weights, inputs, rounding, or intermediate activations can flip outcomes near the decision boundary. Thus, what is feasible is a robust rule: output $1$ when the fraction of $1$’s is clearly above $1/2$ and output 0 when clearly below, with an “uncertain band” around $1/2$.
$\mathrm{AC}^{0}$ already captures approximate majority.
It is important that $\mathrm{AC}^{0}$ can compute approximate majority-type predicates of this form (constant-depth polynomial-size approximate counting/thresholding), while it cannot compute parity (even with nontrivial correlation) [9, 11, 19, 13]. In other words, the practical learnability of majority-like behaviour does not force us to adopt $\mathrm{TC}^{0}$ as the feasible model.
Why majority does not “bootstrap” into parity in practice.
In $\mathrm{TC}^{0}$, one can compose exact threshold/majority gates to compute functions far beyond $\mathrm{AC}^{0}$ (including parity). But this relies on reliability: if each majority gate is only an approximation with even a small error probability, the error typically amplifies under composition, and parity is particularly unforgiving. Parity is maximally sensitive: flipping any single bit flips the output, so small independent errors in intermediate computations destroy correlation with the true parity. Put differently, parity requires a level of sharpness and global consistency that approximate majority does not provide.
Consequences for NIM learnability.
Optimal NIM play is governed by nim-sum, which is global XOR/parity structure (Theorem 3.4). Approximate majority-style computations do not help: even a small error rate in the parity bits of nim-sum produces systematically wrong move choices. Hence it is consistent that:
  • majority-like tasks are learnable/feasible, because coarse thresholds are robust and $\mathrm{AC}^{0}$-compatible;
  • parity-like tasks (and nim-sum) remain infeasible for practical fixed-precision bounded-depth networks;
  • multi-frame representations can restore feasibility by reducing the required parity computations to local parity on a constant number of changed heaps (Lemma 4.5).
This is precisely why an $\mathrm{AC}^{0}$-based model is a useful theoretical abstraction for practical learnability.
5.4 The Multi-Frame Advantage: From Global to Local Parity
While previous architectures, such as the original AlphaGo, utilized multi-frame stacks to capture temporal dynamics [16], we demonstrate that in impartial games, the transition from single-frame to multi-frame representation represents a fundamental shift in required circuit complexity.
  • Single-Frame Obstruction: A network restricted to a single position $P$ must compute the global nim-sum $\bigoplus_{i=1}^{n}x_{i}$. We prove this is equivalent to a PARITY-type computation, which is non-computable in $\mathrm{AC}^{0}$ [9].
  • Multi-Frame Discovery: By providing two frames $(P,P^{\prime})$, the network can isolate the changed heap $i$ and compute the local difference $\Delta=x_{i}\oplus x_{i}^{\prime}$. This local operation is provably within $\mathrm{AC}^{0}$ bounds (Lemma 4.5).
  • Empirical question. Section 6.2 tests whether two-frame policies learn the restoration response substantially more easily than single-frame policies when training data is organised around $(P\text{-position},\text{perturbed})$ transitions. We also discuss teacher-guided MCTS training (Section 6.3) as a more end-to-end but harder-to-bootstrap regime.

### v3

@@ 5.3 Multi-frame possibility results
5.3.1 Two-frame local restoration rule
@@
Suppose the opponent moves from a position $Q$ with $\operatorname{nim}(Q)=0$ to a position $Q^{\prime}$ (so it is now our turn). Let $\Delta(Q,Q^{\prime})=\operatorname{nim}(Q)\oplus\operatorname{nim}(Q^{\prime})$ denote the nimber difference caused by the opponent’s move. If we respond with a move $Q^{\prime}\to R$ such that
@@
By definition, $\operatorname{nim}(Q^{\prime})=\operatorname{nim}(Q)\oplus\Delta(Q,Q^{\prime})$. Since $\operatorname{nim}(Q)=0$ (by hypothesis), we have $\operatorname{nim}(Q^{\prime})=\Delta(Q,Q^{\prime})$. The player’s move creates a new position $R$ with $\operatorname{nim}(R)=\operatorname{nim}(Q^{\prime})\oplus\Delta(Q^{\prime},R)$. Substituting the condition $\Delta(Q^{\prime},R)=\Delta(Q,Q^{\prime})$:
$\operatorname{nim}(R)\;=\;\operatorname{nim}(Q^{\prime})\oplus\operatorname{nim}(Q^{\prime})\;=\;0.$
Thus, matching the opponent’s difference restores the invariant. ∎
Moreover, a concrete restoration move can be selected in $\mathrm{AC}^{0}$ from $(Q,Q^{\prime})$ by computing $\delta=\Delta(Q,Q^{\prime})$ and then choosing (with a fixed tie-break rule) a heap $i$ such that $x_{i}^{\prime}=x_{i}\oplus\delta<x_{i}$ (Appendix C).
[…]
Figure 2: A worked restoration example (three heaps). Starting from a $P$-position $Q$ (green), the opponent changes heap 2 ($5\to 4$), creating a nimber difference $\delta=1$. The agent instantly restores the $P$-position by changing heap 1 ($3\to 2$), which imposes the matching difference $3\oplus 2=1$.
Interpretation (“mirror the delta”).
Once we have made the opponent face a $P$-position, we do not need to recompute $\operatorname{nim}(\cdot)$ from scratch. We only need to mirror the observed change $\Delta(Q,Q^{\prime})$ in our response move. Figure 2 illustrates this mechanism on a concrete three-heap instance.
Proposition 5.5 is structural: it describes a response pattern that keeps returning the game to nim-sum zero, provided the player can (i) compute $\delta=\Delta(Q,Q^{\prime})$ from the last two frames and (ii) choose a legal move $Q^{\prime}\to R$ that realises the same $\delta$. In NIM, when the opponent changes exactly one heap, $\delta$ is simply $x_{i}\oplus x_{i}^{\prime}$, and testing whether a candidate response move realises $\delta$ is a local check on the heap being changed.
@@ 5.3.2 Implementation overview
  • an $\mathrm{AC}^{0}$ selection mechanism for a response move realising a target $\delta$.
5.4 Beyond evaluation: search, rollouts, and the structural gap
Theorem 5.2 shows that no $\mathrm{AC}^{0}$ circuit can evaluate the nim-sum in a single forward pass. A natural follow-up is whether planning—testing candidate moves and simulating play, as in Monte-Carlo Tree Search (MCTS)—can compensate.
We analyse a stylised rollout interface: a planning algorithm queries a rollout policy $\pi$ by simulating a complete game in which both players follow $\pi$ from a start position, and it observes only the terminal win/loss bit. This abstraction isolates the information that a weak policy can expose through repeated simulation; it is not a claim about the optimisation dynamics of AlphaZero-style training (cf. Section 1.2).
Definition 5.6 (Rollout policy and rollout outcome).
A rollout policy is a (deterministic) function $\pi\colon\{0,1\}^{NB}\to\{\text{legal moves}\}$ mapping a NIM position to a legal move. We say $\pi$ is $\mathrm{AC}^{0}$-bounded if it is computed by a circuit of constant depth and size $\mathrm{poly}(NB)$.
A rollout from position $P$ under policy $\pi$ is the terminal play obtained by having both players follow $\pi$ from $P$. The rollout outcome is $\mathrm{RO}_{\pi}(P)\in\{0,1\}$, where $1$ means the first player wins.
Definition 5.7 (Single-policy and multi-policy-head agents).
A single-policy agent $\mathcal{A}=(E,\pi,S)$ consists of an $\mathrm{AC}^{0}$ value head $E$, a single $\mathrm{AC}^{0}$ rollout policy $\pi$, and a search budget $S$ (number of rollout queries). Each rollout query simulates a complete game in which both players follow $\pi$ from a specified start position, returning a single win/loss bit.
A $k$-policy-head agent $\mathcal{A}=(E,\pi_{1},\dots,\pi_{k},S)$ has $k$ distinct $\mathrm{AC}^{0}$ rollout policies. For each of its $S$ rollout queries, the agent selects a policy head $\pi_{j}$ and a start position, then observes the terminal outcome of a complete game played under $\pi_{j}$. The choice of which head to deploy may depend adaptively on the outcomes of previous queries.
Remark 5.8 (Why multiple heads are not redundant).
Since $\mathrm{AC}^{0}$ is closed under parallel composition, one could compute all $k$ policies simultaneously as a single $\mathrm{AC}^{0}$ circuit. The distinction is not about computational power at the circuit level but about the rollout interface: each rollout query simulates a complete game under one policy, and at each game step only that policy determines the next move. A single rollout trajectory under a fixed policy $\pi$ produces one outcome bit; merging $k$ circuits into a single circuit does not change the fact that any one rollout follows a single deterministic rule, yielding a single bit of information. Multiple heads are useful because the agent can issue separate rollout queries under different policies $\pi_{j}$, producing different game trajectories that reveal independent linear functionals of the nim-sum (cf. Propositions 5.9 and 5.10).
Finite-state shootout controller (FSM over policy heads).
Definition 5.7 allows the choice of rollout head to depend adaptively on the outcomes of previous rollout queries. To make this adaptivity explicit (without changing the $\mathrm{AC}^{0}$ constraints on each head), it is convenient to view head-selection as a small finite-state controller that sits outside the heads and only sees the terminal win/loss bit of each rollout.
Formally, fix rollout heads $\pi_{0},\dots,\pi_{k-1}$. A finite-state shootout controller is a tuple
$\mathcal{C}=(\mathcal{S},s_{\mathrm{init}},g,\delta,\mathsf{halt}),$
where $\mathcal{S}$ is a finite set of controller states, $s_{\mathrm{init}}\in\mathcal{S}$ is the initial state, $g\colon\mathcal{S}\to\{0,\dots,k-1\}$ selects which head to use for the next rollout query, $\delta\colon\mathcal{S}\times\{0,1\}\to\mathcal{S}$ updates the controller state after observing a rollout outcome bit $r=\mathrm{RO}_{\pi_{g(s)}}(P)$, and $\mathsf{halt}\subseteq\mathcal{S}$ is a set of halting states associated with an output decision. (When $\mathsf{halt}=\emptyset$, the controller simply runs for a fixed number of queries and outputs a decision at the end.)
This viewpoint does not increase the per-step computational power of the rollout heads: each $\pi_{j}$ remains $\mathrm{AC}^{0}$-bounded. Instead, the controller organises which $\mathrm{AC}^{0}$ head is iterated in a shootout and when to stop querying. Figure 3 depicts a particularly simple controller that turns the $B$ structured rollout heads from Proposition 5.9 into a sequential “bit-query” procedure for deciding whether $\sigma(P)=\operatorname{nim}(P)$ is zero (and hence whether $P$ is a $P$-position).
[…]
Figure 3: A minimal finite-state shootout controller for deciding whether $\sigma(P)=\operatorname{nim}(P)$ is zero using the $B$ rollout heads of Proposition 5.9. The controller queries heads sequentially and halts as soon as it can certify $\sigma(P)\neq 0$. Reaching the final transition to $\sigma(P)=0$ implicitly means $r_{0}=\cdots=r_{B-1}=0$.
Integration (channels vs. protocol). It is helpful to reinterpret Part (ii) of Theorem 5.14 in an “information-channel” language: each policy head $\pi_{j}$ defines a distinct rollout measurement channel which, when iterated to termination, returns a single bit $r_{j}=\mathrm{RO}_{\pi_{j}}(P)$ that depends on the hidden $B$-bit invariant $\sigma(P)$ (cf. the identities in Proposition 5.9). With a single head, repeated rollouts merely resample the same linear functional $\ell(\sigma)$ (Proposition 5.10), so additional search budget cannot reveal the missing bits. With multiple heads, the agent has access to multiple independent channels and can reconstruct $\sigma(P)$ from their outcomes (Proposition 5.9). The finite-state shootout controller introduced above is the natural generalisation of this idea: it does not change what any head can compute, but organises the heads into an adaptive querying protocol—the controller decides which channel to query next (and when to stop, or when to re-query for robustness) based on previously observed rollout outcomes. In this view, multi-head provides the channels, while the FSM provides the sequential control that turns those channels into a practical evaluation procedure.
Depth amplification.
Although a single $\mathrm{AC}^{0}$ circuit cannot compute parity, iterating an $\mathrm{AC}^{0}$ policy for $L$ game steps yields an effective depth $\Theta(L)$ computation, which can reveal parity information. The next proposition makes this explicit by exhibiting $B$ simple $\mathrm{AC}^{0}$ rollout heads whose outcomes recover all $B$ bits of the nim-sum.
Proposition 5.9 (Multi-policy-head rollout sufficiency).
A $B$-policy-head $\mathrm{AC}^{0}$ agent achieves strong mastery of NIM on $N$ heaps of bit-width $B$, using total computation $O(BN^{2}\cdot 4^{B})$. For $B=O(\log N)$ this is polynomial in $N$.
Proof outline.
For each $j\in\{0,\dots,B-1\}$ define a deterministic rollout policy $\pi_{j}$ that repeatedly subtracts $2^{j}$ from the leftmost heap of size at least $2^{j}$ (and otherwise subtracts $1$ from the leftmost nonzero heap). A rollout under $\pi_{j}$ has a well-defined total number of moves whose parity is
$\mathrm{RO}_{\pi_{0}}(P)=\sigma_{0},\qquad\mathrm{RO}_{\pi_{j}}(P)=\sigma_{j}\oplus\sigma_{0}\quad(j\geq 1),$
where $\sigma=\operatorname{nim}(P)$ and $\sigma_{j}$ denotes its $j$th bit. The search controller first queries $\pi_{0}$ to obtain $\sigma_{0}$, then queries each $\pi_{j}$ ($j\geq 1$) to obtain $\sigma_{j}\oplus\sigma_{0}$, and recovers $\sigma_{j}=(\sigma_{j}\oplus\sigma_{0})\oplus\sigma_{0}$; thus $B$ rollouts under $B$ distinct heads recover all $B$ bits of $\sigma$. Testing candidate moves (by running the same $B$ rollouts from each successor position) identifies a move to nim-sum 0. A complete proof is in Appendix F. ∎
A single policy head yields only one bit.
The preceding result relies on having multiple rollout channels. With one fixed rollout head, repeated simulation does not reveal additional independent information about the nim-sum.
Proposition 5.10 (Single-policy information bound).
Fix $j$ and the rollout policy $\pi_{j}$ from the proof outline of Proposition 5.9. For any position $P$, the rollout outcome $\mathrm{RO}_{\pi_{j}}(P)$ is a fixed linear functional $\ell_{j}(\sigma(P))$ of the nim-sum bits (indeed, $\ell_{0}(\sigma)=\sigma_{0}$ and $\ell_{j}(\sigma)=\sigma_{j}\oplus\sigma_{0}$ for $j\geq 1$). Moreover, if $P_{i}$ is obtained from $P$ by modifying one heap (as when evaluating candidate moves), then
$\mathrm{RO}_{\pi_{j}}(P_{i})=\ell_{j}(\sigma(P))\oplus c_{i},$
where the correction bit $c_{i}$ is computable from the known heap values and the chosen modification. Hence all rollout queries under a single head reveal copies of the same unknown bit $\ell_{j}(\sigma(P))$.
Proof.
The linearity claim follows from the explicit formulas above. If $P_{i}$ differs from $P$ by changing one heap from $h$ to $h^{\prime}$, then $\sigma(P_{i})=\sigma(P)\oplus(h\oplus h^{\prime})$ and $\ell_{j}(\sigma(P_{i}))=\ell_{j}(\sigma(P))\oplus\ell_{j}(h\oplus h^{\prime})$, where $\ell_{j}(h\oplus h^{\prime})$ is known from the candidate move. ∎
Remark 5.11 (Exponential signal decay under noisy rollouts).
Proposition 5.10 analyses deterministic rollout heads. In practice, a learned policy is noisy: at each step, it selects the intended move with probability $1-\varepsilon$ and errs otherwise. The terminal outcome of a rollout of length $L$ then depends on the parity of the number of correct moves along the path.
If per-step errors are approximately independent, the standard identity for parity of biased Bernoulli sums gives
$\left|\Pr[\text{first player wins}]-\tfrac{1}{2}\right|\;=\;\tfrac{1}{2}(1-2\varepsilon)^{L},$
which decays exponentially in the rollout length $L$. Even strong per-move accuracy (e.g. $\varepsilon=0.01$) yields a near-fair-coin signal once $L\gtrsim 50$.
Majority voting over $M$ independent rollouts does not remove this bottleneck: distinguishing a bias of $(1-2\varepsilon)^{L}$ from $\tfrac{1}{2}$ requires
$M=\Omega\!\bigl((1-2\varepsilon)^{-2L}\bigr),$
which is exponential in $L$. This quantifies the gap in Figure 4(a): a single noisy rollout head produces not merely limited information but exponentially weak information about the parity-governed outcome, regardless of search budget.
The independence assumption is an idealisation; correlated errors can change constants, but the qualitative message of exponential decay with rollout length remains.
Two-frame rollouts can act as universal verifiers.
The two-frame restoration rule from Section 5.3.1 leads to a rollout policy whose single outcome bit already decides whether a position is a $P$-position.
Definition 5.12 (Universal verifier).
A rollout policy $\pi^{*}$ is a universal verifier if $\mathrm{RO}_{\pi^{*}}(P)=0$ iff $\operatorname{nim}(P)=0$ (equivalently, $P$ is a $P$-position), when both players follow $\pi^{*}$. In other words, the first player loses the rollout precisely when the start position is losing for the player to move.
Proposition 5.13 (Restoration is a universal verifier in the rollout interface).
The two-frame restoration strategy $\pi^{*}$ of Proposition 5.5 is a universal verifier for NIM and is computable in $\mathrm{AC}^{0}$.
Proof.
In the deterministic rollout interface of Definition 5.6, each query is launched from a specified start transition. From that point onward, whenever one player induces a transition with nimber difference $\delta$, the other player can compute $\delta$ from two frames and apply Proposition 5.5 to return to nim-sum 0. Thus the key invariant is that the second player to act after departure from a $P$-position restores nim-sum 0.
If the queried start position satisfies $\operatorname{nim}(P)=0$, the first mover must leave nim-sum nonzero, and the second mover repeatedly restores nim-sum 0; hence the first mover loses and $\mathrm{RO}_{\pi^{*}}(P)=0$. If $\operatorname{nim}(P)\neq 0$, the verifier bit flips correspondingly in this rollout interface, yielding $\mathrm{RO}_{\pi^{*}}(P)=1$. The concrete move-construction details used in Theorem 5.14(iii) are given in Appendix G.
The policy uses only the two-frame difference signal from Lemma 5.4 and fixed tie-breaking, so it is $\mathrm{AC}^{0}$-computable. ∎
Summary theorem.
Combining these ingredients yields a clean separation between single-policy single-frame planning and the two bypasses (multiple policy heads or two-frame inputs). Figure 4 provides a visual summary of this rollout-level separation.
[…]
Figure 4: The Rollout Gap. (a) A single-frame policy acts as a random walk on the parity hypercube; the signal mixes rapidly to $0.5$ (noise). (b) A two-frame agent acts as a Universal Verifier, executing a deterministic restoration path that preserves the truth value of the move perfectly.
Theorem 5.14 (The structural gap in search complexity).
Consider NIM on $N$ heaps of bit-width $B\geq 2$.
  1. (i)
    Single-frame, single-policy limitation. Let $\mathcal{A}=(E,\pi_{j},S)$ be a single-frame $\mathrm{AC}^{0}$ agent using one rollout head $\pi_{j}$ from above. Then all $S$ rollout queries reveal at most one unknown bit $\ell_{j}(\sigma(P))$ of the $B$-bit nim-sum. Consequently, on a random $N$-position the probability that $\mathcal{A}$ identifies a winning move is at most $2^{-(B-1)}+\mathrm{negl}(N)$.
  2. (ii)
    Single-frame, multi-policy-head sufficiency. A $B$-policy-head $\mathrm{AC}^{0}$ agent achieves strong mastery with total computation $O(BN^{2}\cdot 4^{B})$.
  3. (iii)
    Two-frame, single-policy sufficiency. A two-frame $\mathrm{AC}^{0}$ agent using the restoration policy $\pi^{*}$ achieves strong mastery with total computation $O(N^{2}\cdot 4^{B})$ by testing candidate first moves via the universal verifier (accepting those with $\mathrm{RO}_{\pi^{*}}=0$, indicating a $P$-position) and then maintaining $\operatorname{nim}=0$ by restoration.
Proof outline.
Part (i) combines Proposition 5.10 with standard $\mathrm{AC}^{0}$ lower bounds implying negligible correlation with parity-type predicates [7, 24, 18]: since all rollout outcomes under a single head are determined by one unknown parity bit (plus known data), and the agent’s decision function is $\mathrm{AC}^{0}$ in the remaining inputs, the agent cannot distinguish among the $2^{B-1}$ nim-sum values consistent with that bit. Parts (ii) and (iii) follow from Propositions 5.9 and 5.13 respectively. A complete proof is in Appendix G. ∎
@@ 6 Learning impartial games with neural networks
The results above are primarily representational: they identify what information a bounded-depth, fixed-scale quantised model can or cannot compute from its input, and what additional interfaces (history or multiple rollout heads) make optimal play possible. This section connects the theory to the training protocols used in Section 7 and highlights the main modelling choice ($\mathrm{AC}^{0}$ vs. $\mathrm{TC}^{0}$).
6.1 Practical takeaways
  • Single-frame is brittle to parity. In the worst case, fixed-scale quantised bounded-depth inference cannot reliably represent the nim-sum or a winning move selector (Section 5.2).
  • Two bypasses. A two-frame input exposes the transition signal $\Delta(P,P^{\prime})$ and supports restoration (Section 5.3.1), while multiple rollout heads can recover the missing nim-sum bits through depth amplification (Section 5.4).
  • Optimisation remains nontrivial. Even when a solution is expressible in $\mathrm{AC}^{0}$, discovering it from sparse win/loss rewards can be difficult; our experiments therefore isolate a dense restoration supervision signal and separately report search-based performance.
6.2 Teacher-guided data as a controlled source of transitions
Let $\pi_{\mathrm{opt}}$ denote an optimal NIM policy (move to nim-sum 0 whenever possible). For a noise parameter $\varepsilon\in[0,1]$, define a noisy teacher $\pi_{T}^{\varepsilon}$ that plays according to $\pi_{\mathrm{opt}}$ with probability $1-\varepsilon$ and otherwise chooses a uniformly random legal move. Occasional random moves broaden state coverage and, crucially for two-frame agents, generate diverse transition pairs $(Q,Q^{\prime})$ from which the local difference signal $\Delta(Q,Q^{\prime})$ can be inferred.
6.3 Approximate majority, parity, and the $\mathrm{AC}^{0}$ abstraction
One might object that modern networks can implement majority-like decisions, suggesting $\mathrm{TC}^{0}$ (threshold circuits) as the “right” complexity class. Indeed, if one allows scale-adaptive thresholds/biases or fan-in dependent re-scaling, constant-depth threshold models can represent parity-type computations.
Our focus is a different deployment regime: fixed-latency inference with fixed-scale, bounded-range quantisation, where each unit’s threshold is $O(1)$ in units of the smallest nonzero weight. In this regime, each gate has a constant-size witness and collapses to a constant-width DNF, yielding an $\mathrm{AC}^{0}$ abstraction (Theorem 2.4). Constant-depth $\mathrm{AC}^{0}$ circuits can implement approximate counting [1], yet they have exponentially small correlation with parity [5, 7, 24, 18]. This aligns with the empirical difficulty of learning parities [26, 27, 3] and supports using $\mathrm{AC}^{0}$ as a conservative abstraction for fixed-scale quantised inference. Section 6.4 gives a complementary Fourier/spectral interpretation of this local-versus-global parity separation.
6.4 A spectral viewpoint: low-degree Fourier mass vs. parity
A useful complementary lens is Fourier analysis on the Boolean hypercube. Let $f:\{0,1\}^{n}\to\mathbb{R}$ and define the Walsh–Hadamard characters
$\chi_{S}(x)\;=\;(-1)^{\sum_{i\in S}x_{i}},\qquad S\subseteq[n].$
The set $\{\chi_{S}\}_{S\subseteq[n]}$ is an orthonormal basis with respect to the uniform distribution on $\{0,1\}^{n}$. The Fourier coefficients are
$\hat{f}(S)\;=\;\mathbb{E}_{x\sim U(\{0,1\}^{n})}\bigl[f(x)\chi_{S}(x)\bigr],\qquad\text{and}\qquad f(x)=\sum_{S\subseteq[n]}\hat{f}(S)\chi_{S}(x).$
We refer to $|S|$ as the Fourier degree (or “frequency”) of the coefficient $\hat{f}(S)$.
Definition 6.1 (Fourier tail / spectral bandwidth).
For $k\in\mathbb{N}$, define the degree->k>k Fourier tail
$\mathrm{Tail}_{>k}(f)\;=\;\sum_{S\subseteq[n]:\,|S|>k}\hat{f}(S)^{2}.$
We say $f$ is $(k,\varepsilon)$-spectrally band-limited if $\mathrm{Tail}_{>k}(f)\leq\varepsilon$.
Parity is a pure high-frequency character.
For any nonempty $S\subseteq[n]$, the parity function $\mathrm{PARITY}_{S}(x)=\chi_{S}(x)$ has Fourier expansion supported on a single set: $\widehat{\mathrm{PARITY}_{S}}(S)=1$ and $\widehat{\mathrm{PARITY}_{S}}(T)=0$ for $T\neq S$. Thus its spectral “frequency” is exactly $|S|$.
$\mathrm{AC}^{0}$ has little high-frequency mass.
A seminal theorem of Linial–Mansour–Nisan shows that $\mathrm{AC}^{0}$ functions have almost all Fourier mass on low degrees: for $f$ computed by an $\mathrm{AC}^{0}_{d}$ circuit of size $M$, the tail above degree $k$ decays rapidly with $k$ (see [8], and sharper bounds in [25]). In particular, for $M=\mathrm{poly}(n)$ and constant depth $d$, any $\mathrm{AC}^{0}$ function has exponentially small correlation with parity on $\omega(\log^{c}n)$ bits (for fixed $c$).
Local vs. global parity in NIM.
In our encoding, the $j$th bit of the nim-sum is the parity of the $j$th bits across all $N$ heaps, hence a degree-$N$ character. By contrast, the $j$th bit of the two-frame difference $\Delta(P,P^{\prime})$ depends only on the old and new $j$th heap-bit (NIM moves change one heap), hence is a degree-$2$ character. This spectral separation mirrors the single-frame barrier and the two-frame restoration bypass.
Connection to optimisation (informal).
Fourier degree also aligns with a common empirical observation in deep learning: gradient-based training tends to fit low-frequency components before high-frequency ones (“spectral bias” / “frequency principle”) [17, 28]. Since parity is a maximally high-frequency character on the hypercube, this provides a complementary optimisation-based explanation for why parity-type invariants are difficult to learn in practice even when they are representable.
6.5 From global invariants to local checks
The core representational point is that the global invariant $\operatorname{nim}(P)$ is hard to compute from a single frame, whereas the transition quantity $\Delta(P,P^{\prime})$ is locally computable from two frames (Lemma 5.4). Section 7.2 isolates this effect by training one-frame and two-frame models on the same restoration targets.